OpenStudy (anonymous):

A statement Sn about the positive integers is given. Write statements Sk and Sk+1, simplifying Sk+1 completely. Sn: 1 + 4 + 7 + . . . + (3n - 2) = n(3n - 1)/2

OpenStudy (anonymous):

\[Sn: 1 + 4 + 7 + . . . + (3n - 2) = \frac{n(3n - 1)}{2} \] \[Sk: 1 + 4 + 7 + . . . + (3k - 2) = \frac{k(3k - 1)}{2} \]

OpenStudy (anonymous):

\[Sk: 1 + 4 + 7 + . . . + (3k - 2) = \frac{k(3k - 1)}{2}\] \[Sk: 1 + 4 + 7 + . . . + (3(k+1) - 2) = \frac{(k+1)(3(k+1) - 1)}{2}\]

OpenStudy (anonymous):

then algebra to "simplify" the last bit

OpenStudy (anonymous):

for example \[(3(k+1)-2)=3k+3-2=3k+1\]

OpenStudy (anonymous):

and \[\frac{(k+1)(3(k+1) - 1)}{2}=\frac{(k+1)(3k+3) - 1)}{2}=\frac{(k+1)(3k+2)}{2}\]

OpenStudy (bibby):

\(S_{k+1}: 1 + 4 + 7 + . . . + (3(k+1) - 2) = \frac{(k+1)(3(k+1) - 1)}{2}\)?

OpenStudy (anonymous):

I don't get it ? what are the dots ?